# Interact-RAG: Reason and Interact with the Corpus, Beyond Black-Box Retrieval

### v1

@@ 1 Introduction
Large Language Models (LLMs) have shown advancements in natural language understanding and generation but are constrained by their training data, which can be static, outdated, or lack domain-specific knowledge (Huang et al., 2025). Retrieval-Augmented Generation (RAG) has emerged as a prevailing solution to this limitation (Lewis et al., 2020; Gao et al., 2023). By retrieving information from external corpora, RAG systems enable LLMs to access up-to-date information, incorporate specialized knowledge, and reason over proprietary data (Hui et al., 2024; Li et al., 2025c).
The development of RAG has progressed through three stages. The initial approach, Static RAG, performs a single retrieval to fetch relevant documents for the LLM (Gao et al., 2023). To handle more complex tasks, Iterative RAG frameworks were introduced. These systems employ multi-step retrieval pipelines to progressively gather information (Trivedi et al., 2023; Jiang et al., 2023; Chan et al., 2024). The current frontier is Agentic RAG, which uses an LLM-centric agent to autonomously orchestrate the entire workflow with more flexibility (Gao et al., 2025). The agent decides when to retrieve, what to query, and how to analyze the retrieved information (Singh et al., 2025). Advanced implementations include prompt-driven multi-agent workflows (Nguyen et al., 2025; Li et al., 2025b) and other end-to-end trained agents using Supervised Fine-Tuning (SFT) and Reinforcement Learning (RL) to improve reasoning and adaptability (Jin et al., 2025a; Zheng et al., 2025; Qian & Liu, 2025).
Despite these advances, existing agentic RAG frameworks share a critical limitation: they treat the retrieval process as an opaque black-box. The agent is confined to issuing a query and passively receiving text chunks, typically from an embedding-based semantic retriever (Gao et al., 2025; Jin et al., 2025a). This paradigm prevents the agent from inspecting the internal state of the retrieval process, thereby forcing it to relinquish fine-grained control over the process. Consequently, the agent’s exploration is restricted to a trial-and-error loop of query reformulation, which limits the breadth, depth, and overall efficacy of its information seeking. For example, when asked, “Which film was released first, The Jaws of Death or Failure to Launch?”, an agent might first query for the “release date of The Jaws of Death”. This retrieval may fail if the supporting evidence is phrased differently (e.g., “…The Jaws of Death is a 1976 thriller film…”) or if the retriever is distracted by semantically similar but irrelevant entities (e.g., a film named The Hound of Death). Faced with such a failure, existing agents can only resort to repeatedly paraphrasing the query (e.g., “when was The Jaws of Death released”). This often leads to an inefficient loop that fails to obtain the necessary information.
@@
To overcome this limitation, we introduce Interact-RAG, a novel paradigm that transforms the agent from a passive query issuer to an active participant in the retrieval process. Our core idea is to dismantle the retrieval “black box” by providing the agent with transparent and fine-grained control over its information seeking. To achieve this, we propose a lightweight Corpus Interaction Engine, which equips the agent with a versatile set of Interaction Primitives, categorized into three action types: (1) Multi-Faceted Retrieval, which allows the agent to employ diverse retrieval strategies (e.g., semantic, exact) and adaptively fuse their results with different weights; (2) Anchored Matching, which focuses the search on a specific entity to mitigate distraction from irrelevant content; (3) Context Shaping, which enables the agent to proactively manage the retrieval context by retaining efficient documents and adjusting the retrieval scope. As shown in Figure 1, this suite of primitives enables fined-grained control, beyond simple query reformulation.
However, just providing these interactive capabilities is insufficient. Empowering the LLM to strategically master the interactive pipeline remains challenging. First, it is difficult to directly instruct an LLM to manage the intricate multi-step process. To address this, we design a reasoning-enhanced workflow that decomposes the task into three modules: a global planner, an adaptive reasoner, and an executor. This approach not only provides a robust, training-free solution but also synthesizes high-quality agent trajectories for subsequent training. Second, achieving full autonomy requires the model to internalize the strategic policies. Therefore, we leverage the synthesized trajectories and apply Supervised Fine-Tuning (SFT), followed by refinement with Reinforcement Learning (RL). As shown in Figure 1, we finally yield a unified, end-to-end agent capable of executing the entire pipeline, without relying on an explicit multi-module architecture.
@@ 3.1 Interactive Engine and Paradigm
Lightweight Implementation.
Our engine is designed for efficiency and minimal overhead. We implement primitives like exact_search and entity_match leveraging the Full-Text Search (FTS) modules in relational databases (SQLite, 2025), which builds a lightweight text index. The Context Shaping primitives are realized through simple filters. This lightweight approach avoids the overhead of heavy operations or extra LLM invocation. More details are in Appendix C.2.
@@ 3.2 Reasoning-Enhanced Workflow
Directly prompting an LLM to master the entire interactive pipeline is challenging. Therefore, we develop a reasoning-enhanced workflow, decomposing the agent action into a hierarchical and iterative structure. It not only serves as a robust training-free solution but also generates high-quality data to train our end-to-end agent. As shwon in Figure 2, the workflow contains three collaborative modules: a global-planner, an adaptive-reasoner, and an executor.
@@
This modular design clearly decouples high-level planning, detailed reasoning and precise execution. For a general-purpose LLM, this separation is critical, as the well-defined and focused tasks elicit more reliable output. This workflow yields two significant advantages. First, as a training-free method, it enhances the stability and logical coherence of zero-shot RAG. Second, it serves as a data synthesis engine to train the autonomous agent. With logically-structured modules, the LLM operates in a non-reasoning mode to produces clean reasoning traces, free from the verbose and irrelevant thinking content, which is common in native large reasoning models (LRMs).
@@ 3.3 End-to-End Agent Training
Trace Sampling and Fine-Tuning. The initial SFT stage aims to teach the LLM the fundamental mechanics, such as planing, reasoning, and mastering the interactions. We leverage our reasoning-enhanced workflow to generate a collection of trajectories based on QA pairs. To ensure the data quality, we retain only successful trajectories, where the agent’s final answer matches the ground truth. The agent is then fine-tuned on these high-quality trajectories. The training objective is to predict the sequence of thoughts and actions in an auto-regressive manner. During loss calculation, we mask out the tokens of retrieved information, avoiding the distraction during learning.
Policy Refinement with Reinforcement Learning. We then employ RL to enable superior strategies through active exploration. We adopt Group Relative Policy Optimization (GRPO) (Shao et al., 2024), an advanced optimization algorithm, to further refine the agent’s policy $\pi_{\theta}$.
@@
Here, each trajectory incurs an initial penalty of -1. The agent should generate a format-coherent output to overcome this penalty. $\mathbb{I}\{\cdot\}$ denotes the indicator function, which returns 1 if its enclosed condition is true, and 0 otherwise. First, the term $\mathbb{I}\{\tau_{\text{valid}}\}$, grants a +1+1 reward if $\tau$ is syntactically valid, thereby neutralizing the initial penalty. Syntactic validity encompasses the entire action sequence structure, the reasoning format, and the tool call syntax. Second, $\mathbb{I}\{\tau_{\text{valid}}\}\cdot\mathbb{I}\{y_{\text{ans}}\}$ provides a +1+1 reward for task success, where the final answer $y_{\text{ans}}$ matches the ground-truth. This reward is gated by the trajectory’s validity, ensuring that only well-formed output can be rewarded.
@@ 4 Experiments
Datasets. We conduct experiments across six prominent and standard RAG benchmarks. These include two single-hop question-answering datasets, Natural Questions (NQ(Kwiatkowski et al., 2019) and PopQA (Mallen et al., 2023), and four multi-hop question-answering datasets: HotpotQA (Yang et al., 2018), 2WikiMultiHopQA (2Wiki(Ho et al., 2020), MuSiQue (Trivedi et al., 2022), and Bamboogle (Press et al., ). More details are in Appendix C.1
@@
Implementation Details. Following previous works (Jin et al., 2025a; Qian & Liu, 2025), we process the 2018 Wikipedia dump as the retrieval corpus. We employ the E5 (Wang et al., 2022) model as the retriever, fetching the top 3 relevant chunks by default. For all experiments, we use Qwen3-8B (Yang et al., 2025), a recent instruction-tuned model, as the backbone LLM. For training-driven baselines (i.e., Search-R1, S-DeepSearcher, and R-Search), we utilize their official checkpoints trained on Qwen-2.5-7B, since their 8B versions are not available yet. To ensure the comprehensiveness, we also report our results on Qwen2.5-7B in Appendix B.1.
We train the agent on the combined training splits of NQ, HotpotQA, and MuSiQue, and evaluate it on the test splits of all six benchmarks. This setup enables the generalization on both in-distribution and out-of-distribution (PopQA, 2Wiki, Bamboogle). For the training process, we first employed Qwen-Plus to synthesize 4.8K agent trajectories for SFT. Subsequently, we utilized 7.1K question-answer pairs for the RL phase. More details are in Appendix C.3.
@@ 4.1 Main Results
Table 1: Overall performance in Exact Match (EM) and F1 scores across various benchmarks. Bold indicates the best performance, and underline indicates the second-best. The default backbone LLM is Qwen3-8B, while methods marked with a dagger () use their official 7B models. Besides, we also report the results on Qwen2.5-7B in Appendix B.1.
We evaluate Interact-RAG on six benchmarks, with the main results in Table  4. Our findings highlight three key advantages of our approach. First, Interact-RAG consistently achieves best performance across all datasets. On average, it improves the EM-score by 9.7 points (22.5% relative gain) and the F1-score by 9.5 points (18.0% relative gain) over the second-best method, Search-R1. Second, the performance gains are more pronounced on complex multi-hop QA tasks. For instance, on Musique, Interact-RAG delivers a 36.4% relative improvement. Concurrently, it maintains strong performance on single-hop benchmarks like NQ and PopQA, with relative improvements of 11.0% and 12.3% on EM scores. This validates the effectiveness of our interaction-reasoning paradigm in tackling complex challenges. Third, our trained agent demonstrates great generalization. Trained with train-splits of HotpotQA, Musique, and NQ, it achieves consistent improvements on both in-distribution and out-of-distribution benchmarks. This indicates that the learned capability are not task-specific, underscoring the robustness and generalizability of our approach.
@@ 4.2 Ablation Study
Impact of the Training Strategy. For our two-phase training, removing SFT leads to severe performance drops, especially on challenging datasets like Musique (-21.9%). This highlights its role in building fundamental mechanics of planing, reasoning, and iterative interaction. Similarly, omitting RL also causes marked declines, as RL is essential to develop more strategic policies. These results demonstrate that while SFT establishes the core patterns of reasoning and interaction, RL further optimizes the agent’s policy to achieve better performance. (More discussion in Section 4.4).
@@ 4.3 Training-Free Scenarios
In scenarios with limited training resources or requiring zero-shot deployment, training-free solutions are practically important. Therefore, we evaluate our training-free approach, termed Interact-RAG-Workflow. As shown in Table 3, our approach consistently outperforms MA-RAG across various benchmarks, underscoring the intrinsic effectiveness of our reasoning-interaction paradigm even without model training. To better understand the impact of individual components, we conduct two ablation studies. First, removing the interaction (i.e, resorting to a black-box query-search) leads to a significant performance drop, highlighting the critical role of fine-grained retrieval control. Second, “w/o workflow” means omit our reasoning-enhanced workflow and directly instruct the LLM through an end-to-end prompt (detailed in Appendix C.4). This also results in performance degradation, confirming the effectiveness of our workflow to orchestrate the entire RAG process.
@@ 4.4 Detailed Analysis
Figure 3: Comparison of the efficiency of information retrieval.
@@
Efficiency of Information Retrieval. We assess the retrieval efficiency by measuring the number of action iterations. We compare our Interact-RAG against two query-only methods: an ablation variant restricted to only the query-search action (termed as Ours-Search) and the Search-R1 method. The results in Figure 4 indicate that Interact-RAG always achieves the highest EM scores with the minimum action iterations. This efficiency is particularly pronounced on complex multi-hop datasets (2Wiki and Musique), where tasks demand more intricate information seeking. This finding validates the core advantage of our paradigm: by providing the agent with fine-grained control, it can navigate the information space more effectively, avoiding inefficient trial-and-error loops. (A case study is in Figure 6).

### v3

@@ 1 Introduction
Large Language Models (LLMs) have shown advancements in natural language understanding and generation but are constrained by their training data, which can be static, outdated, or lack domain-specific knowledge (Huang et al., 2025). Retrieval-Augmented Generation (RAG) has emerged as a prevailing solution to this limitation (Lewis et al., 2020; Gao et al., 2023). By retrieving information from external corpora, RAG systems enable LLMs to access up-to-date information, incorporate specialized knowledge, and reason over proprietary data (Hui et al., 2024; Li et al., 2025f).
The development of RAG has progressed through three stages. The initial approach, Static RAG, performs a single retrieval to fetch relevant documents for the LLM (Gao et al., 2023). To handle more complex tasks, Iterative RAG frameworks were introduced. These systems employ multi-step retrieval pipelines to progressively gather information (Trivedi et al., 2023; Jiang et al., 2023; Chan et al., 2024). The current frontier is Agentic RAG, which uses an LLM-centric agent to autonomously orchestrate the entire workflow with more flexibility (Gao et al., 2025; Li et al., 2025c). The agent decides when to retrieve, what to query, and how to analyze the retrieved information (Singh et al., 2025; Chen et al., 2025). Advanced implementations include prompt-driven multi-agent workflows (Nguyen et al., 2025; Li et al., 2025b) and other end-to-end trained agents using Supervised Fine-Tuning (SFT) and Reinforcement Learning (RL) to improve reasoning and adaptability (Jin et al., 2025a; Zheng et al., 2025; Qian and Liu, 2025).
Despite these advances, existing agentic RAG frameworks share a critical limitation: they treat the retrieval process as an opaque black-box. The agent is confined to issuing a query and passively receiving text chunks, typically from an embedding-based semantic retriever (Gao et al., 2025; Jin et al., 2025a). This paradigm prevents the agent from inspecting the internal state of the retrieval process, thereby forcing it to relinquish fine-grained control over the process. Consequently, the agent’s exploration is restricted to a trial-and-error loop of query reformulation, which limits the breadth, depth, and overall efficacy of its information seeking. For example, when asked, “Which film was released first, The Jaws of Death or Failure to Launch?”, an agent might first query for the “release date of The Jaws of Death”. This retrieval may fail if the supporting evidence is phrased differently (e.g., “…The Jaws of Death is a 1976 thriller film…”) or if the retriever is distracted by semantically similar but irrelevant entities (e.g., a film named The Hound of Death). Faced with such a failure, existing agents can only resort to repeatedly paraphrasing the query (e.g., “when was The Jaws of Death released”). This often leads to an inefficient loop of blind guessing that fails to obtain the necessary information.
@@
To overcome this limitation, we introduce Interact-RAG, a novel paradigm that transforms the agent from a passive query issuer to an active participant in the retrieval process. Our core idea is to dismantle the retrieval “black box” by providing the agent with transparent and fine-grained control over its information seeking. To achieve this, firstly, we propose a Corpus Interaction Engine, which equips the agent with a versatile set of Interaction Primitives, categorized into three action types: (1) Multi-Faceted Retrieval, which allows the agent to employ diverse retrieval strategies (e.g., semantic, exact) and adaptively fuse their results with different weights; (2) Anchored Matching, which focuses the search on a specific entity to mitigate distraction from irrelevant content; (3) Context Shaping, which enables the agent to proactively manage the retrieval context by retaining efficient documents and adjusting the retrieval scope. This suite of primitives serves as a foundation for the fined-grained control, beyond simple query reformulation (as shown in in Figure 1).
However, just providing these interactive capabilities is insufficient. Empowering the LLM to actively and strategically master the interactive pipeline remains challenging. First, it is difficult to directly instruct an LLM to manage the intricate multi-step process. To address this, we design a reasoning-enhanced workflow that decomposes the task into three modules: a global planner, an adaptive reasoner, and an executor. This approach not only provides a robust, training-free solution but also synthesizes high-quality agent trajectories for subsequent training. Second, achieving full autonomy requires the model to internalize the strategic policies. Therefore, we leverage the synthesized trajectories and apply Supervised Fine-Tuning (SFT), followed by refinement with Reinforcement Learning (RL). As shown in Figure 1, we finally yield a unified, end-to-end agent capable of executing the entire pipeline, without relying on an explicit multi-module architecture.
@@ 3.1 Interactive Engine and Paradigm
Implementation.
Our engine is designed for computational efficiency with small overhead. We implement primitives like exact_search and entity_match by leveraging the Full-Text Search (FTS) modules in relational databases (SQLite, 2025). While this builds an additional text index, the computational cost is negligible. Furthermore, the Context Shaping primitives are implemented through simple filters. It is worth noting that while the agent may invoke multiple strategies in a single iteration, our engine avoids generating multiple large contexts. Instead, it produces a single consolidated context by aggregating these primitives. Further implementation details and validation are provided in Appendix D.2.
@@ 3.2 Reasoning-Enhanced Workflow
Directly prompting an LLM to master the entire interactive pipeline is challenging. Therefore, we develop a reasoning-enhanced workflow, decomposing the agent action into a hierarchical and iterative structure. It not only serves as a robust training-free solution but also generates high-quality data to train our end-to-end agent. As shown in Figure 2, the workflow contains three collaborative modules: a global-planner, an adaptive-reasoner, and an executor.
@@
This modular design clearly decouples high-level planning, detailed reasoning and precise execution. For a general-purpose LLM, this separation is critical, as the well-defined and focused tasks elicit more reliable output. This workflow yields two significant advantages. First, as a training-free method, it enhances the stability and logical coherence of zero-shot RAG. Second, it serves as a data synthesis engine to train the autonomous agent. With logically-structured modules, the LLM operates in a non-reasoning mode to produce clean reasoning traces, free from the verbose and irrelevant thinking content, which is common in native large reasoning models (LRMs).
@@ 3.3 End-to-End Agent Training
Trace Sampling and Fine-Tuning. The initial SFT stage aims to teach the LLM the fundamental mechanics, such as planning, reasoning, and mastering the interactions. We leverage our reasoning-enhanced workflow to generate a collection of trajectories based on QA pairs. To ensure the data quality, we retain only successful trajectories, where the agent’s final answer matches the ground truth. The agent is then fine-tuned on these high-quality trajectories. The training objective is to predict the sequence of thoughts and actions in an auto-regressive manner. During loss calculation, we mask out the tokens of retrieved information, avoiding the distraction during learning.
Policy Refinement with Reinforcement Learning. We then employ RL to enable superior strategies through active exploration. We adopt Group Relative Policy Optimization (GRPO) (Shao et al., 2024), an advanced optimization algorithm, to further refine the agent’s policy $\pi_{\theta}$. During the policy updating, we also mask out the tokens of retrieved information.
@@
Here, each trajectory incurs an initial penalty of -1. The agent should generate a format-coherent output to overcome this penalty. $\mathbb{I}\{\cdot\}$ denotes the indicator function, which returns 1 if its enclosed condition is true, and 0 otherwise. First, the term $\mathbb{I}\{\tau_{\text{valid}}\}$, grants a +1+1 reward if $\tau$ is syntactically valid, thereby neutralizing the initial penalty. Syntactic validity encompasses the entire action sequence structure, the reasoning format, and the tool call syntax. Second, $\mathbb{I}\{\tau_{\text{valid}}\}\cdot\mathbb{I}\{y_{\text{ans}}\}$ provides a +1+1 reward for task success, where the final answer $y_{\text{ans}}$ matches the ground-truth. This reward is gated by the trajectory’s validity, ensuring that only well-formed output can be rewarded. It is worth noting that, the inclusion of the “-1 initial penalty” was primarily intended for intuitive logic. Mathematically, this constant term is canceled out during the group-based normalization process in GRPO and does not influence the advantages or the actual training.
@@ 4 Experiments
4.1 Experimental Settings
Datasets. We conduct experiments across six prominent and standard RAG benchmarks. These include two single-hop question-answering datasets, Natural Questions (NQ(Kwiatkowski et al., 2019) and PopQA (Mallen et al., 2023), and four multi-hop question-answering datasets: HotpotQA (Yang et al., 2018), 2WikiMultiHopQA (2Wiki(Ho et al., 2020), MuSiQue (Trivedi et al., 2022), and Bamboogle (Press et al., ). More dataset details are in Appendix D.1 Besides, for more comprehensive evaluation over non-wiki domains, we conduct additional experiments on the MultiHop-RAG benchmark (Tang and Yang, 2024) (more details in Appendix C.2).
@@ 4.1 Experimental Settings
Experimental Details. Following previous works (Jin et al., 2025a; Qian and Liu, 2025), we process the 2018 Wikipedia dump as the retrieval corpus. We employ the e5-base-v2 (Wang et al., 2022) model as the retriever, fetching the top 3 relevant chunks. To ensure a fair comparison, we maintain the same corpus and retriever across all methods, and re-evaluate all baselines under this unified setting. For all experiments, we use Qwen3-8B by default (Yang et al., 2025), a recent instruction-tuned model. For training-driven baselines (i.e., Search-R1, S-DeepSearcher, and R-Search), we utilize their official checkpoints trained on Qwen-2.5-7B, since their 8B versions are not available yet. To ensure the comprehensiveness, we also report our results on Qwen2.5-7B in main results.
We train our agent on the combined training splits of NQ, HotpotQA, and MuSiQue, and evaluate it on the test splits of all six benchmarks. This setup enables the generalization on both in-distribution and out-of-distribution (PopQA, 2Wiki, Bamboogle). For the training process, we first employed Qwen-Plus to synthesize 4.8K agent trajectories for SFT. Subsequently, we utilized 7.1K question-answer pairs for the RL phase. More details are in Appendix D.3.
@@ 4.2 Main Results
Results on Qwen2.5-7B Backbone
Interact-RAG-7B 47.8 61.6 63.6 71.0 30.9 39.5 47.6 61.1 43.7 52.9 51.6 54.4 47.5 56.8
Results on Qwen3-8B Backbone
@@
Table 1: Overall performance in Exact Match (EM) and F1 scores across various benchmarks. Bold and underline denote the best and second-best performance. For the Qwen3-8B results, methods marked with a dagger () use their official 7B models, due to the lack of 8B versions. Notably, our Interact-RAG was trained on 12K QA data, a small fraction of the 170K QA pairs used for Search-R1. This data disparity also explains Search-R1’s higher results on the NQ dataset when using the 7B model.
We evaluate Interact-RAG on six benchmarks, with the main results in Table  LABEL:tab:qa_results. Our findings highlight three key advantages of our approach. First, Interact-RAG consistently achieves best performance across all datasets. On average, it improves the EM-score by 9.7 points (22.5% relative gain) over the second-best method, Search-R1. And our 7B version also achieves a relative improvement of 10.1%. Notably, our Interact-RAG was trained on 12K QA data, a small fraction of the 170K QA pairs used for Search-R1. This data disparity also explains Search-R1’s higher results on the NQ dataset using the 7B model. Second, the performance gains are more pronounced on complex multi-hop QA tasks. For instance, on Musique, Interact-RAG delivers a 36.4% relative improvement. Concurrently, it maintains strong performance on single-hop benchmarks like NQ and PopQA, with relative improvements of 11.0% and 12.3% on EM scores. This validates the effectiveness of our interaction-reasoning paradigm in tackling complex challenges. Third, our trained agent demonstrates great generalization. Trained with train-splits of HotpotQA, Musique, and NQ, it achieves consistent improvements on both in-distribution and out-of-distribution benchmarks. This indicates that the learned capability are not task-specific, underscoring the robustness and generalizability of our approach.
@@ 4.3 Ablation Study
Impact of the Training Strategy. For our two-phase training, removing SFT leads to severe performance drops, especially on challenging datasets like Musique (-21.9%). This highlights its role in building fundamental mechanics of planning, reasoning, and iterative interaction. Similarly, omitting RL also causes marked declines, as RL is essential to develop more strategic policies. These results demonstrate that while SFT establishes the core patterns of reasoning and interaction, RL further optimizes the agent’s policy to achieve better performance. (More discussion in Section 4.5).
Ablation on Interaction Primitives. To isolate the specific contributions of each component, we conduct a fine-grained ablation study on the interaction primitives (Table 3). The results reveal a strong synergistic effect: while removing the single module degrades performance, the complete absence of interaction leads to the most significant drop. This confirms that these primitives function collectively to surpass black-box retrieval. Furthermore, the contributions of different components vary across data patterns, showing the comprehensiveness of our design. For example, Multi-Faceted Retrieval and Anchored Matching are vital for 2Wiki, which may demand explicit facts. And Context Shaping proves more impactful on Musique, where the prevalence of distractors requires robust context scaling.
Method 2Wiki Musique PopQA
Interact-RAG 69.6 34.8 56.0
   w/o All-Interaction 63.4 30.1 50.2
   w/o Multi-faceted Retrieval 66.0 34.6 55.1
   w/o Anchored Matching 66.3 34.4 53.4
   w/o Context Shaping 68.8 33.6 55.2
Table 3: Fine-grained ablation study on interaction primitives, reported in Exact Match (EM) scores.
@@ 4.4 Training-Free Scenarios
In scenarios with limited training resources or requiring zero-shot deployment, training-free solutions are practically important. Therefore, we evaluate our training-free approach, termed Interact-RAG-Workflow. As shown in Table 4, our approach consistently outperforms MA-RAG across various benchmarks, underscoring the intrinsic effectiveness of our reasoning-interaction paradigm even without model training. To better understand the impact of individual components, we conduct two ablation studies. First, removing the interaction (i.e, resorting to a black-box query-search) leads to a significant performance drop, highlighting the critical role of fine-grained retrieval control. Second, “w/o workflow” means omit our reasoning-enhanced workflow and directly instruct the LLM through an end-to-end prompt (detailed in Appendix D.4). This also results in performance degradation, confirming the effectiveness of our workflow to orchestrate the entire RAG process.
@@ 4.5 Detailed Analysis
Figure 3: Comparison of the retrieval iterations and the accuracy.
@@
Efficiency of Information Retrieval. We further assess the retrieval process by measuring the number of action iterations. We compare our Interact-RAG against two query-only methods: an ablation variant restricted to only the query-search action (termed as Ours-Search) and the Search-R1 method. The results in Figure 4 indicate that Interact-RAG always achieves the highest EM scores with the minimum action iterations. This is particularly pronounced on complex multi-hop datasets (2Wiki and Musique), where tasks demand more intricate information seeking. This finding validates the core advantage of our paradigm: by providing the agent with fine-grained control, it can navigate the information space in less iterations, avoiding inefficient trial-and-error loops. (A case study is in Figure 6).